# A novel hybrid time-varying graph neural network for traffic flow forecasting

### v1

@@ V-F2 Effect analysis of different types of graph structure in TVGCRM and HTVGNN
HTVGNN-tg: The time-varying mask embedding is excluded from the RTP multi-head attention, while TVGCRM remains unchanged as HTVGNN.
HTVGNN-wc: Remove the coupling part of the time-varying graph.
HTVGNN-sc: Simultaneously eliminate RTP multi-head self-attention and coupled time-varying graphs in a synchronized manner.
HTVGNN: This is a hybrid time-varying graph convolutional network,as shown in Fig. 1
The results presented in Table 7 demonstrate the superiority of HTVGNN over both HTVGNN-tg and HTVGNN-wc, highlighting the positive impact of incorporating time-varying mask embedding on enhancing self-attention accuracy. Furthermore, the superior performance of HTVGNN compared to HTVGNN-wc indicates its advantage in capturing spatial correlations in transportation networks by leveraging coupling time-varying graphs.
In particular, The ablation study revealed that our proposed HTVGNN, incorporating time-varying graphs, yields significant benefits in reducing the cumulative prediction error. Detailed comparison results are presented in Table III.
The results presented in Table III unequivocally demonstrate the superiority of HTVGNN with coupling time-varying graphs over its counterpart without such coupling. This advantage becomes increasingly pronounced as the time step increases. Furthermore, Table 6 reveals that HTVGNN outperforms HTVGNN-wc by 3.92%, 4.04%, and 4.32% in terms of MAE, RMSE, and MAPE on PEMS04 dataset at a temporal resolution of 60 minutes, respectively. Similarly, on PEMS08 dataset at the same temporal resolution, HTVGNN achieves improvements of 5.43%, 3.20%, and 4.57% for MAE, RMSE, and MAPE.
@@ V-G Hyperparameters analysis of HTVGNN
We analyzed the time-varying graph embedding dimension represented by E, the model embedding dimension represented by D, the number of ETP multi-head self-attention heads represented by Heads, and the number of layers of the model represented by Layers. These four hyperparameters have significant effects on PEMS04 and PEMS08 data. The influence of the set of experimental results, the value of D on the data set PEMS04 is 4 to 10, and the value of D on the data set PEMS08 is 2 to 8.
@@ V-G1 Sensitivity to Layers
The impact of the model’s layer count is illustrated in Fig. 5a. However, our findings indicate that increasing the number of layers does not yield a significant improvement in performance. This observation suggests that the HTVGNN model may have reached its predictive error limit.
@@ V-G2 Sensitivity to Heads
The impact of the number of heads on RTP multi-head attention is illustrated in Fig. 5b. It can be observed from the figure that while the attention score can be effectively computed with two heads, the dimensionality of the corresponding space for this head count is insufficient. As the number of heads increases, spatio-temporal information is mapped to a higher-dimensional space, resulting in more reasonable correlation coefficient scores. However, an excessively large number of heads may introduce numerous negative high-dimensional spaces, which hinders accurate calculation of the attention score.
@@ V-G3 Sensitivity to Dimensions
The impact of the hidden embedding dimension D is illustrated in Fig. 5c. As depicted, smaller dimensions fail to effectively encode spatiotemporal information, while excessively large ones may lead to overfitting.
@@ V-G4 Sensitivity to E
The impact of the size E in the embedding space for time-varying graphs is illustrated in Fig. 5d. From the figure, it can be observed that a smaller dimension E fails to capture spatial correlations effectively, while an excessively large dimension E leads to overfitting and suboptimal learning of time-varying graphs.
@@ V-H Visualiation
In order to visually demonstrate the superiority of our proposed model, we present results of various benchmark models at 500 time steps on PEMS04 and PEMS08 datasets. The comparison between traffic flow prediction results and actual values is depicted in Fig.7. It can be observed that the HTVGNN model exhibits enhanced accuracy and responsiveness towards abnormal traffic flow, while also demonstrating reduced error in handling complex and dynamic traffic patterns.
@@ V-I Efficiency Study
Additionally, we conducted a comparison of the training and inference efficiency of HTVGNN, Bi-STAT, and PDFormer on the PEMS04 dataset. To ensure fairness in our evaluation, all methods were executed on a server equipped with an Intel Core i7 13700KF processor and a single NVIDIA RTX graphics card. The results of this comparison are presented in Table 9.
As shown in Table 9, despite the time-consuming nature of RNNs, the HTVGNN model utilizes a single layer. In contrast, PDFormer and Bi-STAT solely rely on transformer-based architectures that inherently support parallel processing. However, to capture deep spatio-temporal correlations, transformer-based models often require multiple layers; for instance, PDFormer employs six layers while Bi-STAT uses two layers. Nevertheless, due to the inclusion of a recollection process, this model incurs twice the original time complexity. Consequently, these transformer-only models do not offer any advantage in terms of time complexity. Conversely, our proposed HTVGNN model achieves superior prediction error performance on PEMS04 dataset with just one layer despite employing computationally intensive RNNs. Moreover, HTVGNN significantly outperforms PDFormer and Bi-STAT in prediction accuracy while maintaining similar computational requirements.
@@ VI Conclusion
In this paper, we propose a novel hybrid time-varying graph neural network prediction method. To accurately capture the most relevant temporal features in temporal correlation modeling, we introduce a time-aware multi-head attention mechanism based on time-varying mask enhancement. We allocate the optimal attention calculation scheme for different temporal features. In spatial correlation modeling, we employ two strategies. Firstly, in static spatial correlation modeling, we design a coupled graph learning mechanism to integrate the graph learned at each time step and effectively enhance the model’s long-term prediction capability. Secondly, in dynamic spatial correlation modeling within the road network, we define a mask based on topological matrix and semantic matrix to optimize the dynamic time-varying graph and improve the model’s ability to capture dynamic spatial correlations. Simulation experiments conducted on four real datasets validate our proposed method HTVGNN by demonstrating its superior prediction accuracy compared with state-of-the-art spatio-temporal graph neural network models. Furthermore, our proposed coupling graph learning method significantly improves long-term prediction performance of the model as observed in ablation experiments. Future work will focus on exploring more effective masking techniques through model pre-training and providing more efficient yet powerful masking mechanisms for Transformer models.

### v3

@@ V-F2 Effect analysis of different types of graph structure in CTVGCRM and HTVGNN
: A novel hybrid time-varying graph neural network,as shown in Fig. 1
: The time-varying mask embedding is excluded from the RTP multi-head attention, while CTVGCRM remains unchanged as HTVGNN.
: Remove the coupling part of the time-varying graph.
: The time-varying mask embedding is excluded from the RTP multi-head attention, and the coupling part of the time-varying graph is removed.
The results presented in Table VIII demonstrate the superiority of HTVGNN over both w/o-tm, highlighting the positive impact of incorporating time-varying mask embedding on enhancing self-attention accuracy. Furthermore, the superior performance of HTVGNN compared to w/o-cg and the superior performance of w/o-cg compared to w/o-bc indicates its advantage in capturing spatial correlations in transportation networks by leveraging coupling time-varying graphs.
In particular, The ablation study revealed that our proposed HTVGNN, incorporating time-varying graphs, yields significant benefits in reducing the cumulative prediction error. Detailed comparison results are presented in Table IX.
The results presented in Table IX unequivocally demonstrate the superiority of HTVGNN with coupling time-varying graphs over its counterpart without such coupling. This advantage becomes increasingly pronounced as the time step increases. Furthermore, Table IX reveals that HTVGNN outperforms HTVGNN-wc by 3.92%, 4.04%, and 4.32% in terms of MAE, RMSE, and MAPE on PEMS04 dataset at a temporal resolution of 60 minutes, respectively. Similarly, on PEMS08 dataset at the same temporal resolution, HTVGNN achieves improvements of 5.43%, 3.20%, and 4.57% for MAE, RMSE, and MAPE.
@@ V-G Hyperparameters analysis of HTVGNN
We analyzed the time-varying graph embedding dimension represented by E, the model embedding dimension represented by D, the number of enhanced temporal perception multi-head self-attention heads represented by Heads, and the number of layers of the model represented by Layers. These four hyperparameters have significant effects on PEMS08 data. The influence of the set of experimental results, ie, the value of D on the data set PEMS08, is 2 to 10.
@@ V-G1 Sensitivity to Layers
The impact of the model’s layer count is illustrated in Figure 6(a). However, our findings indicate that increasing the number of layers does not yield a significant improvement in performance. This observation suggests that the HTVGNN model may have reached its predictive error limit.
@@ V-G2 Sensitivity to Heads
The impact of the number of heads on RTP multi-head attention is illustrated in Figure 6(b). It can be observed from the figure that while the attention score can be effectively computed with two heads, the dimensionality of the corresponding space for this head count is insufficient. As the number of heads increases, spatio-temporal information is mapped to a higher-dimensional space, resulting in more reasonable the attention score. However, an excessively large number of heads may introduce numerous negative high-dimensional spaces, which hinders accurate calculation of the attention score.
@@ V-G3 Sensitivity to Dimensions
The impact of the hidden embedding dimension D on the model HTVGNN is illustrated in Figure 6(c). As depicted, smaller dimensions fail to effectively encode spatiotemporal information, while excessively large ones may lead to overfitting.
@@ V-G4 Sensitivity to E
The impact of the size E in the embedding space for time-varying graphs is illustrated in Figure 6(d). From the figure, it can be observed that a smaller dimension fails to capture spatial correlations effectively, while an excessively large dimension E leads to overfitting and suboptimal learning of time-varying graphs.
@@ V-H Visualiation
In order to visually demonstrate the superiority of our proposed model, we present the traffic flow of various benchmark models over time steps on PEMS04 and PEMS08 datasets. The comparison between traffic flow prediction results and actual values is depicted in Figure 7. It can be observed that the HTVGNN model exhibits enhanced accuracy and responsiveness towards abnormal traffic flow, while also demonstrating reduced error in handling complex and dynamic traffic changes.
@@ V-I Efficiency Study
Additionally, we conducted a comparison of the training and inference efficiency of HTVGNN, Bi-STAT, and PDFormer on the PEMS08 dataset. To ensure fairness in our evaluation, all methods were executed on a server equipped with an Intel Core i7 13700KF processor and a single NVIDIA RTX3090 graphics card. The results of this comparison are presented in Table X.
| Model | Traning(sec/epochs) | Inference(sec/samples) |
| PDFormer | 56 | 0.0025 |
| Bi-STAT | 61 | 0.0024 |
| HTVGNN | 50 | 0.0061 |
As shown in Table X, despite the time-consuming nature of RNNs, the HTVGNN model utilizes a single layer. In contrast, PDFormer and Bi-STAT solely rely on transformer-based architectures that inherently support parallel processing. However, to capture deep spatio-temporal correlations, transformer-based models often require multiple layers; for instance, PDFormer employs six layers while Bi-STAT uses two layers. Nevertheless, due to the inclusion of a recollection process, this model incurs twice the original time complexity. Consequently, these transformer-only models do not offer any advantage in terms of time complexity. Conversely, our proposed HTVGNN model achieves superior prediction error performance on PEMS08 dataset with just one layer despite employing computationally intensive RNNs. Moreover, HTVGNN significantly outperforms PDFormer and Bi-STAT in prediction accuracy while maintaining similar computational requirements.
@@ VI Conclusion
In this paper, we propose a novel hybrid time-varying graph neural network prediction method, HTVGNN. To accurately capture the most relevant temporal features in temporal correlation modeling, we introduce a time-aware multi-head attention mechanism based on time-varying mask enhancement. We allocate the optimal attention calculation scheme for different temporal features. In spatial correlation modeling, we employ two strategies. Firstly, in static spatial correlation modeling, we design a coupled graph learning mechanism to integrate the graph learned at each time step and effectively enhance the model’s long-term prediction capability. Secondly, in dynamic spatial correlation modeling within the road network, we define a mask based on topological matrix and semantic matrix to optimize the dynamic time-varying graph and improve the model’s ability to capture dynamic spatial correlations. Simulation experiments conducted on four real datasets validate our proposed method HTVGNN by demonstrating its superior prediction accuracy compared with state-of-the-art spatio-temporal graph neural network models. Furthermore, our proposed coupling graph learning method significantly improves long-term prediction performance of the model as observed in ablation experiments. Future work will focus on exploring more effective masking techniques through model pre-training and providing more efficient yet powerful masking mechanisms for Transformer models.